OpenStudy (anonymous):

the length of a side of a rectangle is 2 unit less than a number, and the width is 4 units more than the length. which of the following expressions represents area of the rectangle, in square units?

OpenStudy (jiteshmeghwal9):

let the unknown no. be 'x'. so, length=(x-2) width=(x-2)+4 area of a rectangle=width * length so, (x-2) *(x-2+4)=area

OpenStudy (anonymous):

thank you